# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v3

@@ VIII.1.4 Comment M4
Further to Comment M3, the distributions of the individual matrix elements of a general $2\times 2$ random matrix $\mathcal{M}$ are not fundamentally important for determining the eigenvalue spacing distribution of $\mathcal{M}$. According to Eqs. (10) and (87), the distribution of the mixture of random variables that comprise the quantity $\mathcal{D}(\mathcal{M})$ determines the spacing distribution. Thus, the individual matrix elements need not consist of sums of different power-transformed exponential random variables; it is sufficient that the discriminant $\mathcal{D}(\mathcal{M})$ be proportional to the square of a Weibull random variable with shape parameter $\tau=q+1$:
$\mathcal{D}(\mathcal{M})=kL^{2},~{}~{}k\neq 0,~{}L\sim\text{Weibull}(\theta=\sigma(\sigma_{1},\sigma_{2},\ldots,\sigma_{l}),\tau=q+1),$ (113)
@@
where $\theta=\sigma(\cdots)$ denotes the fact that the Weibull scale parameter $\theta$ can depend on any distributional scale parameters $\{\sigma_{1},\sigma_{2},\ldots,\sigma_{l}:l\in\mathbb{N}\}$ the random-matrix model may involve.
@@
$S=\left(U^{2}+V^{2}\right)^{1\over{q+1}}\sim\text{Weibull}\left(\theta=2\sigma^{2}_{R},\tau=q+1\right),$ (118)
and that the distribution of $Z\equiv{S/\mu_{S}}$ (i.e., the mean-scaled spacing distribution) is again the Brody distribution. Comparing (12) and (118), it is then clear that model $\mathcal{M}_{R}$ satisfies condition (8) (with $k=1$) only in the special case when $\sigma_{R}=\sqrt{\sigma_{\text{e}}/2}$. Therefore, a key point is that the above example satisfies condition (113) (with $k=1$ and $L=R^{2\over{q+1}}$) but does not (in general) satisfy condition (8). The reason is that condition (8) [with $f_{Y}(y)$ as given by Eq. (7)] and condition (113) are not identical; the former is a special case of the latter.
@@
There is a fundamental distinction between conditions (8) and (113): condition (8) is a sufficiency condition for model (6) to have a Brody spacing distribution, whereas condition (113) is a more general sufficiency condition that can be applied to evaluate and construct other types of $2\times 2$ models. The important general conclusion is that Eq. (17) holds whenever condition (113) is satisfied regardless of the distributions of the individual matrix elements. Interestingly, this means that many different $2\times 2$ random-matrix models will have a Brody spacing distribution. The question of uniqueness has therefore been answered in the negative.
VIII.2 Physics Comments
VIII.2.1 Comment P1
In defining model (6), no specific conditions were imposed that demanded the matrices be real and/or symmetric. The interesting consequence is that complex non-Hermitian $2\times 2$ random matrices can have spacing statistics intermediate between Poisson and Wigner. The real-symmetric structure of $2\times 2$ GOE matrices is not required. GOE matrices owe their real and symmetric structure to assumptions and conditions that originate from quantum mechanics. The fact that GOE matrices are symmetric, for example, originally stems from the general assumption that quantum Hamiltonians are Hermitian. This assumption is useful in quantum mechanics since it ensures that all eigenvalues are real, but it is not actually required. In other words, hermiticity is a sufficient but not necessary condition for the eigenvalues to be real. In sub-class A (for example), the eigenvalues are guaranteed to be real by imposition of condition (31d) when the model constants are real or conditions (79) when the constants are complex; hermiticity is not required. Hermiticity is not fundamentally important in the present context due to the small size of the matrices which allows for other tractable (and less restrictive) conditions that can be imposed to ensure the eigenvalues are real; when dealing with large matrices, on the other hand, hermiticity is more imperative. In summary, model (6) has no built-in quantum-mechanical assumptions or conditions.
VIII.2.2 Comment P2
The $q=0$ limit of model (6) can (subject to condition (8)) generate a variety of new classes of $2\times 2$ (correlated) random-matrix models having Poisson spacing statistics. For example, the eigenvalue spacings of real asymmetric random matrices of the (pruned) form (c.f., Sec. IV.2.2)
will have a Poisson distribution as long as the real constants $\{c_{j}:j=1,\ldots,8\}$ are such that condition (8) is satisfied. [^12] Sufficient conditions on $\{c_{j}\}$ are: (i) $c_{2}=c_{8}$; (ii) $c_{4}c_{6}=0$; (iii) $04c_{3}c_{5}-2c_{1}c_{7}=$ 0; (iv) $c^{2}_{1}+4(c_{4}c_{5}+c_{3}c_{6})=0$. One solution (from an infinite number) is: $\{c_{1}=-2,~{}c_{3}=c_{5}=1,~{}c_{2}=c_{4}=c_{8}=0,~{}c_{6}=c_{7}=-1\}$. Valid sub-classes of models generated from the $q=0$ limit of model (6) are interesting in the respect that they illustrate that the matrix elements need not be IID nor do any elements need to be Poisson (exponentially) distributed in order for the eigenvalues to have Poissonian spacings. In the above example, none of the matrix elements are exponentially distributed (except the lower diagonal element when $c_{7}>0$ and $c_{8}=0$). In general, the matrices need not even be real or Hermitian as the $q=0$ limits of models (82) and (85) illustrate. Complex eigenvalues can as well have Poisson spacing statistics (c.f., Sec. V.2). The interesting point is that (real or complex) random matrices whose elements are fully correlated can generate (real or complex) eigenvalues whose spacings are completely uncorrelated. Non-trivial $2\times 2$ random-matrix models having Poisson spacing statistics are curiously absent in the literature.131313“Non-trivial” in this context means $2\times 2$ random-matrix models where all matrix elements are non-zero, independent, and not exponentially distributed. Interestingly, there does however exist a simple $N\times N$ diagonal model whose eigenvalue spacing distribution converges to Eq. (1) as $N\to\infty$ (see Ref. [69] and references therein).
VIII.2.3 Comment P3
The $q=1$ limit of model (6) can (subject to condition (8)) generate many new classes of $2\times 2$ (correlated) random-matrix models having Wigner spacing statistics. For example, the non-Hermitian matrices (from sub-class A)
$\mathcal{M}^{(1)}_{A}(q=1)=\left(\begin{array}[]{cc}X+c_{1}\sqrt{Y}&c_{2}{1\over\sqrt{Y}}\\ c_{3}\sqrt{Y^{3}}&X+c_{4}\sqrt{Y}\end{array}\right)~{}~{}$ (124)
will (under the condition that $k_{A}\equiv(c_{1}-c_{4})^{2}+4c_{2}c_{3}$ is real and non-zero) have Wigner spacing distributions. The existence of such models reinforces the important point that the independence and quantum-mechanical assumptions associated with the construction of the GOE are sufficient but not necessary conditions for the spacing distributions of $2\times 2$ random matrices to be Wignerian. In other words, the elements of a $2\times 2$ random matrix need not be IID nor does the matrix need to be real and/or symmetric in order for its eigenvalues to have a Wigner spacing distribution. A separate and physically important example that illustrates this point can be found in Ref. [70] (see also Ref. [71] for an example of a real asymmetric $2\times 2$ random-matrix model having a Wigner spacing distribution). [^14] Note that, in Refs. [70, 71], the matrix elements are Gaussian distributed. An important point that is often glossed over in the literature is that when the classical independence assumption is relaxed (as is the case in model (6)), the matrix elements need not be Gaussian distributed nor are they required to be in order for the eigenvalue spacings to be Wigner distributed. The preceding is true even when independence is stipulated provided the matrix elements are allowed to be non-identically distributed.
@@ VIII.2.3 Comment P3
VIII.2.4 Comment P4
Pruned and cropped models may seem unphysical due to the (conditionally) fully-correlated nature of the matrix elements (c.f., Comment M2). Random matrix models having highly-correlated entries are actually not a novelty in physics (see, for example, Refs. [74, 75, 76, 77, 78] and references contained therein). The success of classical RMT (in physics and other fields) does not rely on all of its specific model assumptions. As mentioned in the Introduction, there is a large body of mathematical research that rigorously shows that the classical results of RMT do not depend on the specific nature of the matrix elements (including whether or not the entries are correlated) but rather only on the symmetry class of the matrix (e.g., real symmetric, complex Hermitian, etc.). This is the celebrated spectral universality that was originally conjectured by Wigner, Dyson, Gaudin, and Mehta [11]. The classical independence assumption used in the construction of the GOE is an assumption that simplifies the mathematical analysis and allows for analytical results but has no direct connection to quantum mechanics since the matrix elements of quantum Hamiltonian systems are typically not independent (see, for example, Refs. [79, 80]). The matter of whether the matrix elements are uncorrelated, partially correlated, or fully correlated, is therefore not a general concern.
It is worth re-iterating that the full correlation (conditionally) present in pruned and cropped models can (if desired) be selectively destroyed by employing random variables in the summand exponents and/or coefficients (c.f., Comment M2) and/or making use of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$ in the general model (6) (c.f., the paragraph preceding Sec. IV.1). For reference, Table 1 lists the models given in the current paper that possess independent matrix elements. In general, mutual independence between any pair of matrix elements can be effected by judicious placement of random functions. For instance, it is clear from inspection of Table 1 that no example was given that possessed mutually independent off-diagonal elements. This particular independence structure can be easily effected in many ways using random functions. A simple example is:
@@ VIII.2.4 Comment P4
where the random variable $V$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{V}(v)$.
IX Conclusion
@@
Real matrices that derive from model (6) are not necessarily symmetric. The quantum hermiticity condition, which ensures that all eigenvalues are real, was not imposed in defining model (6) and is (in the present $2\times 2$ context) a sufficient but not necessary condition for the eigenvalues to be real. As Examples 2-4 in Sec. VII explicitly demonstrate, complex non-Hermitian $2\times 2$ random matrices whose eigenvalues are real (or complex) can possess spacing statistics intermediate between Poisson and Wigner. The author is not aware of any other complex non-Hermitian random-matrix models (or desymmetrized non-Hermitian physical systems) whose real or complex eigenvalues possess such statistics. It is important to emphasize that the classical assumption of IID matrix elements has been dropped in the present model and the question of whether the preceding statement about non-Hermitian matrices holds when the IID assumption is not removed remains open. There are also questions related to pseudo-symmetry [72] and/or pseudo-hermiticity [73] that are pertinent and interesting but such questions are beyond the scope of the present paper.
Finally, as discussed in Comment M4, the individual matrix elements need not be composed of power-transformed exponential random variables. Thus, there is some flexibility in assigning the distributions of the matrix elements. The immediate question then is whether other models having more tunable degrees of correlation (i.e., more flexibility with the assignment of independence between matrix elements) can be constructed. This is indeed possible. The author hopes to introduce a simple but versatile class of partially-correlated $2\times 2$ random-matrix models in a sequel paper.
@@ A.1 The Weibull Distribution
Let the random variable $W\sim\text{Weibull}(\theta,\tau)$, that is, let $W$ have a Weibull distribution with scale parameter $\theta>0$ and shape parameter $\tau>0$. The p.d.f. of $W$ is then given by:
$f_{W}(w;\theta,\tau)={\tau\over\theta}w^{\tau-1}\exp\left(-{w^{\tau}\over\theta}\right),~{}w>0.$ (128)
@@
$\mu_{W}(\theta,\tau)=\left({1\over\theta}\right)^{-{1/\tau}}\Gamma\left(1+{1\over\tau}\right).$ (129)
@@ Appendix B Derivation of Distribution (107)
The derivation of distribution (107) closely parallels the derivation of the Brody distribution given in Sec. III. The task in this appendix is to determine the distribution of the random spacing $S$ as given by Eq. (11) when $Y\sim\text{Gamma}(\sigma_{\text{g}},\xi=2)$ (instead of $Y\sim\text{Exp}(\sigma_{\text{e}})$). This can be efficiently accomplished by exploiting property (135), which immediately yields:
@@
$f_{V}(v)={(q+1)\over\sigma^{2}_{\text{g}}}v^{2q+1}\exp\left(-{v^{q+1}\over\sigma_{\text{g}}}\right).$ (137)
@@
where $k>0$ is the discriminant constant found in condition (8). Note that the parametric dependence of $f_{V}$ has been suppressed for notational brevity in Eqs. (137)-(138).
The distribution of the standardized or mean-scaled eigenvalue spacing $Z\equiv{S/\mu_{S}}$ can then be calculated as in Sec. III using the transformation equation (16), where in the present case,

### v4

@@ X.1.5 Comment M5
The distributions of the individual matrix elements of a general $2\times 2$ random matrix $\mathcal{M}$ are not fundamentally important for determining the eigenvalue spacing distribution of $\mathcal{M}$. According to Eqs. (10) and (116), the distribution of the mixture of random variables that comprise the quantity $\mathcal{D}(\mathcal{M})$ determines the spacing distribution. Thus, the individual matrix elements need not consist of sums of different power-transformed exponential random variables; it is sufficient that the discriminant $\mathcal{D}(\mathcal{M})$ be proportional to the square of a Weibull random variable with shape parameter $\tau=q+1$:
$\mathcal{D}(\mathcal{M})=kL^{2},~~k\neq 0,~L\sim\text{Weibull}(\kappa=\sigma(\sigma_{1},\sigma_{2},\ldots,\sigma_{l}),\tau=q+1),$ (173)
@@
where $\kappa=\sigma(\cdots)$ denotes the fact that the Weibull scale parameter $\kappa$ can depend on any distributional scale parameters $\{\sigma_{1},\sigma_{2},\ldots,\sigma_{l}:l\in\mathbb{N}\}$ the random-matrix model may involve.
@@
$S=\left(U^{2}+V^{2}\right)^{1\over{q+1}}\sim\text{Weibull}\left(\kappa=2\sigma^{2}_{R},\tau=q+1\right),$ (178)
and that the distribution of $Z\equiv{S/\mu_{S}}$ (i.e., the mean-scaled spacing distribution) is again the Brody distribution.
@@
Condition (8) [with $f_{Y}(y)$ as given by Eq. (7)] and condition (173) are not identical; the former is a particular case of the latter. There is a fundamental distinction between conditions (8) and (173): condition (8) is a sufficiency condition for model (6) to have a Brody spacing distribution, whereas condition (173) is a more general sufficiency condition that can be applied to evaluate and construct other types of $2\times 2$ models. The important general conclusion is that Eq. (17) holds whenever condition (173) is satisfied regardless of the distributions of the individual matrix elements. Interestingly, this means that many different $2\times 2$ random-matrix models will have a Brody spacing distribution. The question of uniqueness has therefore been answered in the negative.
@@
X.1.6 Comment M6
As previously mentioned, the full correlation (conditionally) present in pruned and cropped models can (if desired) be selectively destroyed by employing random variables in the summand exponents and/or coefficients (c.f., Comment M2) and/or making use of the functions $\{g_{1}(X),g_{2}(X),h(T)\}$ in the general model (6) (c.f., the paragraph preceding Section V.1). For reference, Table 1 lists the models given in the current paper that possess independent matrix elements. In general, mutual independence between any pair of matrix elements can be effected by judicious placement of random functions. For instance, it is clear from inspection of Table 1 that no example was given that possessed mutually independent off-diagonal elements. This particular independence structure can be easily effected in many ways using random functions. A simple example is:
@@ X.1.6 Comment M6
where the random variable $V$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{V}(v)$. The above example derives from applying the generalization discussed in Section VIII.2 to sub-class A, more specifically, allowing some of the sub-class A coefficients $\{c_{j}\}$ to be random variables as follows: $\{c_{1},c_{2},c_{3},c_{4}\}=\left\{1,{1\over 2}\sin^{2}(2qV),{1\over 2},2\sin^{2}(qV)\right\}$.
X.2 Physics Comments
X.2.1 Comment P1
In defining model (6), no specific conditions were imposed that demanded the matrices be real and/or symmetric. The interesting consequence is that complex non-Hermitian $2\times 2$ random matrices can have spacing statistics intermediate between Poisson and Wigner. The real-symmetric structure of $2\times 2$ GOE matrices is not required. GOE matrices owe their real and symmetric structure to assumptions and conditions that originate from quantum mechanics. The fact that GOE matrices are symmetric, for example, originally stems from the general assumption that quantum Hamiltonians are Hermitian. This assumption is useful in quantum mechanics since it ensures that all eigenvalues are real, but it is not required. In other words, hermiticity is a sufficient but not necessary condition for the eigenvalues to be real. In sub-class A (for example), the eigenvalues are guaranteed to be real by imposition of condition (31d) when the model constants are real or conditions (108) when the constants are complex; hermiticity is not required. Hermiticity is not fundamentally important in the present context due to the small size of the matrices which allows for other tractable (and less restrictive) conditions that can be imposed to ensure the eigenvalues are real; when dealing with large matrices, on the other hand, hermiticity is more imperative. In summary, model (6) has no built-in quantum-mechanical assumptions or conditions.
X.2.2 Comment P2
The $q=0$ limit of model (6) can (subject to condition (8)) serve as a new class of $2\times 2$ (correlated) random-matrix models having Poisson spacing statistics. Valid sub-classes of models generated from the $q=0$ limit of model (6) are interesting in the respect that they illustrate that the matrix elements need not be IID nor do any elements need to be Poisson (exponentially) distributed in order for the eigenvalues to have Poissonian spacings. For example, the eigenvalue spacings of real non-symmetric random matrices of the (pruned) form (c.f., Section V.2.2)
will have a Poisson distribution as long as the real constants $\{c_{j}:j=1,\ldots,8\}$ are such that condition (8) is satisfied. [^12] Sufficient conditions on $\{c_{j}\}$ are: (i) $c_{2}=c_{8}$; (ii) $c_{4}c_{6}=0$; (iii) $4c_{3}c_{5}-2c_{1}c_{7}=0$; (iv) $c^{2}_{1}+4(c_{4}c_{5}+c_{3}c_{6})=0$. One solution (from an infinite number) is: $\{c_{1}=-2,~c_{3}=c_{5}=1,~c_{2}=c_{4}=c_{8}=0,~c_{6}=c_{7}=-1\}$. Note that none of the matrix elements of model (184) are exponentially distributed (except the lower diagonal element when $c_{7}>0$ and $c_{8}=0$). In general, the matrices need not even be real or Hermitian as the $q=0$ limits of models (111) and (114) illustrate. Complex eigenvalues can as well have Poisson spacing statistics (c.f., Section VII.2). The interesting point is that (real or complex) random matrices whose elements are fully correlated can generate (real or complex) eigenvalues whose spacings are completely uncorrelated. Non-trivial $2\times 2$ random-matrix models having Poisson spacing statistics are curiously absent in the literature.131313“Non-trivial” in this context means $2\times 2$ random-matrix models where all matrix elements are non-zero, independent, and not exponentially distributed. Interestingly, there does however exist a simple $N\times N$ diagonal model whose eigenvalue spacing distribution converges to Eq. (1) as $N\to\infty$ (see Ref. [69] and references therein).
X.2.3 Comment P3
The $q=1$ limit of model (6) can (subject to condition (8)) serve as a new class of $2\times 2$ (correlated) random-matrix models having Wigner spacing statistics. The existence of this class of models reinforces the important point that the independence and quantum-mechanical assumptions associated with the construction of the GOE are sufficient but not necessary conditions for the spacing distributions of $2\times 2$ random matrices to be Wignerian. More explicitly, the elements of a $2\times 2$ random matrix need not be IID nor does the matrix need to be real and/or symmetric in order for its eigenvalues to have a Wigner spacing distribution. As a simple but concrete example, the (possibly complex) sub-class of matrices (c.f., Eq. (32d))
$\mathcal{M}^{(1)}_{A}(q=1)=\left(\begin{array}[]{cc}X+c_{1}\sqrt{Y}&c_{2}{1\over\sqrt{Y}}\\ c_{3}\sqrt{Y^{3}}&X+c_{4}\sqrt{Y}\end{array}\right)$ (187)
will (under the condition that $k_{A}\equiv(c_{1}-c_{4})^{2}+4c_{2}c_{3}$ is real and non-zero) have Wigner spacing distributions. A separate and physically important example that illustrates this point can be found in Ref. [70] (see also Ref. [71] for an example of a real non-symmetric $2\times 2$ random-matrix model having a Wigner spacing distribution). [^14] Note that, in Refs. [70, 71], the matrix elements are Gaussian distributed. An important point that is often glossed over in the literature is that when the classical independence assumption is relaxed (as is the case in model (6)), the matrix elements need not be Gaussian distributed nor are they required to be in order for the eigenvalue spacings to be Wigner distributed. The preceding is true even when independence is stipulated provided the matrix elements are allowed to be non-identically distributed.
X.2.4 Comment P4
The real symmetric models in Section VI may seem unphysical due to the fully-correlated nature of the matrix elements. Random matrix models having highly-correlated entries are actually not a novelty in physics (see, for example, Refs. [74, 75, 76, 77, 78] and references contained therein). The success of classical RMT (in physics and other fields) does not rely on all of its specific model assumptions. As mentioned in the Introduction, there is a large body of mathematical research that rigorously shows that the classical results of RMT do not depend on the specific nature of the matrix elements (including whether or not the entries are correlated) but rather only on the symmetry class of the matrix (e.g., real symmetric, complex Hermitian, etc.). This is the celebrated spectral universality that was originally conjectured by Wigner, Dyson, Gaudin, and Mehta [11]. The classical independence assumption used in the construction of the GOE is an assumption that simplifies the mathematical analysis and allows for analytical results but has no direct connection to quantum mechanics since the matrix elements of quantum Hamiltonian systems are typically not independent (see, for example, Refs. [79, 80]). The matter of whether the matrix elements are uncorrelated, partially correlated, or fully correlated, is therefore not a general concern.
X.2.5 Comment P5
Although model (6) can generate real-symmetric and complex-Hermitian sub-classes of models, it is most suitable for generating real non-symmetric and complex non-Hermitian models. In terms of application, model (6) with arbitrary $f_{Y}(y)$ and bounded $q\geq 0$, could (broadly speaking) serve as a general one-parameter random-matrix model for two-state non-Hermitian (classical or quantum) Hamiltonian systems. Many interesting physical examples of such two-state systems can be found in the literature (see, for example, Refs. [81, 82, 83, 84] and references therein). Model (6) could therefore find application in many different non-Hermitian contexts and settings where the two-state approximation is useful. Specific examples could include (to name only two): the statistics of avoided resonance crossings in open chaotic billiards (e.g., [85]) and the eigenvalue spacing statistics of dielectric microcavities with different refractive indices (e.g., [86]).
XI Conclusion
@@
Real matrices that derive from model (6) are not necessarily symmetric. The quantum hermiticity condition, which ensures that all eigenvalues are real, was not imposed in defining model (6) and is (in the present $2\times 2$ context) a sufficient but not necessary condition for the eigenvalues to be real. As Examples 2-4 in Section IX explicitly demonstrate, complex non-Hermitian $2\times 2$ random matrices whose eigenvalues are real (or complex) can possess spacing statistics intermediate between Poisson and Wigner. The author is not aware of any other complex non-Hermitian random-matrix models (or desymmetrized non-Hermitian physical systems) whose real or complex eigenvalues possess such statistics. It is important to emphasize that the classical assumption of IID matrix elements has been dropped in the present model and the question of whether the preceding statement about non-Hermitian matrices holds when the IID assumption is not removed remains open. There are also questions related to pseudo-symmetry [72] and/or pseudo-hermiticity [73] that are pertinent and interesting but such questions are beyond the scope of the present paper.
Finally, as discussed in Comment M5, the individual matrix elements need not be composed of power-transformed exponential random variables. Thus, there is some flexibility in assigning the distributions of the matrix elements. The immediate question then is whether other models having more tunable degrees of correlation (i.e., more flexibility with the assignment of independence between matrix elements) can be constructed. This is indeed possible. The author hopes to introduce a simple but versatile class of semi-correlated $2\times 2$ random-matrix models in a sequel paper.
@@ A.1 The Weibull Distribution
Let the random variable $W\sim\text{Weibull}(\kappa,\tau)$, that is, let $W$ have a Weibull distribution with scale parameter $\kappa>0$ and shape parameter $\tau>0$. The p.d.f. of $W$ is then given by:
$f_{W}(w;\kappa,\tau)={\tau\over\kappa}w^{\tau-1}\exp\left(-{w^{\tau}\over\kappa}\right),~w>0.$ (188)
@@
$\mu_{W}(\kappa,\tau)=\left({1\over\kappa}\right)^{-{1/\tau}}\Gamma\left(1+{1\over\tau}\right).$ (189)
@@ Appendix B Derivation of Distribution (136)
The derivation of distribution (136) closely parallels the derivation of the Brody distribution given in Section III. The task in this appendix is to determine the distribution of the random spacing $S$ as given by Eq. (11) when $Y\sim\text{Gamma}(\sigma_{\text{g}},\xi=2)$ (instead of $Y\sim\text{Exp}(\sigma_{\text{e}})$). This can be efficiently accomplished by exploiting property (195), which immediately yields:
@@
$f_{V}(v;\sigma_{\text{g}},q)={(q+1)\over\sigma^{2}_{\text{g}}}v^{2q+1}\exp\left(-{v^{q+1}\over\sigma_{\text{g}}}\right).$ (197)
@@
where $k>0$ is the discriminant constant found in condition (8). Note that parametric dependences have been suppressed (for notational brevity) in the first line of Eq. (198).
The distribution of the standardized or mean-scaled eigenvalue spacing $Z\equiv{S/\mu_{S}}$ can then be calculated as in Section III using the transformation equation (16), where in the present case,